a very large umber of balls are thrown vertically upwards in quick succession in such a way that the next ball is thrown when the previous one at maximum height. If the maximum height is 5m the number of balls thrown per minute is????

Grade:11

1 Answers

Arun
4 years ago

Max. Height reached by each ball (h) = 5 m.

Final velocity, v =0

Let Initial velocity=u m/s

=> v^2 = u^2 - 2gh

=> 0=u^2 - 2×10×5

=> u^2 = 100

=> u = 10 m/s

Now, put u=10 in first equation of motion to get the value of time,

V = u - gt

=>0 = 10 -10t

=> t= 1 sec.

Time taken by each ball to reach the max. height = 1 sec.

Conversely,

No. of balls thrown in 1 sec. = 1

No. of balls thrown in 1 min. or 60 sec. =60 balls.